# In Search of Trees: Decision-Tree Policy Synthesis for Black-Box Systems via Search

### v1

@@ 6.3 Experiment #3: Number of Predicate Nodes
The number of predicate nodes directly influences the size of the search space. We study the runtime by considering trees of depth three and varying the maximum number of nodes from three to six.
@@ 7 Further Discussion and Limitations
Our approach is effective at computing small and optimal decision-tree policies. There is an exponential runtime dependency with respect to the size of the tree and the discretisation of the state space. It may be infeasible with our approach to construct large tree policies or deal with high-dimensional environments. It is also possible that not all environments may be controlled by small decision trees.
However when it is applicable, we believe small trees are valuable for interpretability reasons, and our approach provides the means to easily obtain such trees. The exponential runtime factor in our approach is inherent to every approach that aims to provide guarantees.
@@
The optimality is important since it guarantees that we obtain the best performing tree under consideration, which may be relevant for some applications. It also allows us to conclude in cases when no such tree exists, and in general understand the limits of decision trees as control policies.
@@ 8 Conclusion
We presented a novel search-based method for computing an optimal decision-tree policy given a set of initial states and a black-box system. To the best of our knowledge, our approach is the first to consider such a setting. The key component is our trace-based pruning technique, which discards large portions of the search space at runtime. We illustrated the practicality of the approach on classical control benchmarks. When the environment is controllable by a small tree, our approach provides a way to obtain a small and optimal tree despite only requiring black-box access to the system.
@@ 9 Acknowledgements
This research received support from NWO Veni grant Explainable Monitoring (222.119), Independent Research Fund Denmark under reference number 10.46540/3120-00041B, and the Villum Investigator Grant S4OS under reference number 37819. This work was done in part while Anna Lukina was visiting the Simons Institute for the Theory of Computing.
@@ Appendix A Appendix / supplemental material
We considered three environments as defined in Gymnasium [^2] https://github.com/Farama-Foundation/Gymnasium. The dynamics may be found in their git repository, however, our algorithm does not have access to the internal dynamics, and only observes the state-action outputs in a black-box fashion. In the following, we describe the initial states and specifications. Note that specifications are also treated as black-box for the algorithm.
A.1 CartPole

### v2

@@ 6.3 Experiment #3: Number of Predicate Nodes
The number of predicate nodes influences the search space size. We study the runtime for trees of depth three and varying the maximum number of nodes from three to six.
@@ 7 Further Discussion and Limitations
Our approach is effective at computing small and optimal decision-tree policies given a discretisation of the predicates and a set of initial states. There is an exponential runtime dependency with respect to the size of the tree and the discretisation of the state space. It may be infeasible with our approach to construct large tree policies or deal with high-dimensional environments. It is also possible that not all environments may be controlled by small decision trees.
However, when it is applicable, we believe small trees are valuable for interpretability reasons, and our approach provides the means to easily obtain such trees. The exponential runtime factor in our approach is inherent to every approach that aims to provide guarantees.
@@
The optimality is important since it guarantees that we obtain the best performing tree given our definition, which may be relevant for some applications. It also allows us to conclude in cases when no such tree exists, and in general understand the limits of decision trees as control policies.
@@ 8 Conclusion
We presented a novel search-based method for computing an optimal decision-tree policy given a black-box deterministic system, a set of initial states, and a discretisation of the tree predicates. To the best of our knowledge, our approach is the first to consider such a setting. The key component is our trace-based pruning technique, which discards large portions of the search space at runtime. We illustrated the practicality of the approach on classical control benchmarks. When the environment is controllable by a small tree, our approach provides a way to obtain a small and optimal tree despite only requiring black-box access to the system.
@@ 9 Acknowledgements
This research was funded by the NWO Veni grant Explainable Monitoring (222.119) and by the Independent Research Fund Denmark under reference number 10.46540/3120-00041B. This work was done in part while Anna Lukina was visiting the Simons Institute for the Theory of Computing.
@@ Appendix A Appendix / supplemental material
A.1 Proof of Correctness for Trace-Based Pruning
We introduced trace-based pruning in Sections 5.2 and 5.3. Here we provide a proof sketch of the idea.
Reminder. For a given backtracking variable, predicates are explored in increasing order of the threshold. For example, after the predicate $[s_{1}\geq v]$ has been assigned to a backtracking variable, after backtracking, any predicate that would be considered as the next predicate for the backtracking variable must have a greater threshold value, i.e., have the form $[s_{1}\geq v^{\prime}]$ with $v^{\prime}>v$.
In principle, we could explore arbitrary orderings of predicates for a given backtracking variable. For correctness, any ordering would do, as long as we explore all predicates. However, for our trace-based pruning to be effective, it makes book-keeping simpler if we order the predicates by increasing threshold value.
The order of backtracking variables is fixed in our algorithm and does not change during the execution. The variables are indexed by this order, i.e., all combinations for variable $b_{i+1}$ are exhausted before taking the next value for backtracking variable $b_{i}$. The distance $d_{i}$ for backtracking variable $b_{i}$ with corresponding predicate $[s_{j}\geq v]$ is the minimum value of the state dimension $s_{j}$ for which the predicate was evaluated to true amongst all traces that were considered since the predicate $[s_{j}\geq v]$ has been assigned.
Our pruning mechanism is designed for the case when considering all possible trees of a fixed shape, i.e., the positions of the predicate and leaf nodes are fixed, and the goal is to enumerate all such trees. Recall that in our algorithm, we enumerate all tree shapes, and for each shape, we enumerate all corresponding trees where we make use of our trace-based pruning.
Claim. Let us consider a single initial state. Let the backtracking variable $b_{i}$ have a predicate $[s_{j}\geq v]$ assigned with distance value $d_{i}$. When all predicates for all backtracking variables $b_{k}$ with $k>i$ have been exhausted, the goal is to assign the next predicate to backtracking variable $b_{i}$. The claim is that any predicate of the form $[s_{j}\geq v^{\prime}]$ with $v^{\prime}<d_{i}$ will not result in a tree policy that is better than the current best tree policy. In order words, the algorithm should assign a predicate $[s_{j}\geq v^{\prime\prime}]$ with $v^{\prime\prime}=\min\{v:v>d_{i}\}$, or consider predicates with the next state dimension if no such value exists, or further backtrack if that is also not possible.
Proof sketch. To see why trace-based pruning is correct, we give a proof by contradiction. Assume that there is a tree policy $\pi^{*}$ that is better than the current best known policy $\pi^{+}$, and that $\pi^{*}$ has all nodes associated with backtracking variables $b_{m}$ with $m<i$ fixed to the same values as above, the node associated with the backtracking variable $b_{i}$ has a predicate $[s_{j}\geq v^{\prime\prime}]$ with $v^{\prime\prime}<d_{i}$, and the remaining nodes are assigned any predicates.
This means that the trace produced by $\pi^{*}$ has at least one state such that, when querying $\pi^{*}$ for the next action, the tree evaluates the predicate associated with the backtracking variable $b_{i}$ and $s_{j}<d_{i}$. However, this is not possible: since the search explored all possible traces obtainable with the allowed predicates, by definition of $d_{i}$ there cannot be a trace for which $\pi^{*}$ queries the node $b_{i}$ with $s_{j}<d_{i}$. This concludes the proof by contradiction.
A.2 Environments
We considered three environments as defined in the Gymnasium [^2] https://github.com/Farama-Foundation/Gymnasium. The dynamics may be found in their git repository, however, our algorithm does not have access to the internal dynamics, and only observes the state-action outputs in a black-box fashion. In the following, we describe the initial states and specifications. Note that specifications are also treated as black-box for the algorithm.
Reward function.
Our goal is to maximise or minimise the trace length, i.e., the time it takes to reach the goal. Note that this is different from typical reinforcement learning, where the reward function generates a value at each state, and the objective is to minimise or maximise the cumulative reward. In this sense, that reward function is relatively smooth, whereas in our work, the reward is essentially binary: either the goal has been reached or not.
CartPole.
@@ A.2 Environments
The goal is to maximise the trace size, up to a maximum of $10{,}000$ steps (other works usually consider up to $200$ steps).